# Zero-Shot Learning of Causal Models

### v2

@@ Appendix F Comparing Cond-FiP with CausalNF
Appendix G Evaluating Generalization of Cond-Fip to Larger Sample Size
In the main tables (Table 1, Table 2, and Table 3), we evaluated Cond-FiP’s generalization capabilities to larger graphs ($d=50$, $d=100$) than those used for training ($d=20$). In this section, we carry a similar experiment where instead of increasing the total nodes in the graph, we test Cond-FiP on datasets with more samples $n_{\textnormal{test}}=1000$, while Cond-FiP was only trained for datasets with sample size $n_{\textnormal{train}}=400$.
@@ Appendix G Evaluating Generalization of Cond-Fip to Larger Sample Size
Table 27: Results for Noise Prediction with Larger Sample Size ($n_{\textnormal{test}}=1000$). We compare Cond-FiP against the baselines for the task of predicting noise variables from the input observations. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario.
@@
Table 28: Results for Sample Generation with Larger Sample Size ($n_{\textnormal{test}}=1000$). We compare Cond-FiP against the baselines for the task of generating samples from the input noise variables. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario.
@@
Table 29: Results for Interventional Generation with Larger Sample Size ($n_{\textnormal{test}}=1000$). We compare Cond-FiP against the baselines for the task of generating interventional data from the input noise variables. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario.

### v3

@@ Appendix I Experiments on Real World Benchmark
Appendix K Limitations of Cond-FiP
K.1 Evaluating Generalization of Cond-Fip to Larger Sample Size
In the main results (Table 1, Table 2, and Table 3), we evaluated Cond-FiP’s generalization capabilities to larger graphs ($d=50$, $d=100$) than those used for training ($d=20$). In this section, we carry a similar experiment where instead of increasing the total nodes in the graph, we test Cond-FiP on datasets with more samples $n_{\mathcal{D_{\mathrm{test}}}}=1000$, while Cond-FiP was only trained for datasets with sample size $n_{\mathcal{D}}=400$.
@@
Table 27: Results for Noise Prediction with Larger Sample Size ($n_{\mathcal{D_{\mathrm{test}}}}=1000$). We compare Cond-FiP against the baselines for the task of predicting noise variables from the input observations. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Results indicate that Cond-FiP does not yet benefit from larger context sizes at inference, suggesting the need to scale both the model and training data for richer contexts.
@@
Table 28: Results for Sample Generation with Larger Sample Size ($n_{\mathcal{D_{\mathrm{test}}}}=1000$). We compare Cond-FiP against the baselines for the task of generating samples from the input noise variables. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Results indicate that Cond-FiP does not yet benefit from larger context sizes at inference, suggesting the need to scale both the model and training data for richer contexts.
@@
Table 29: Results for Interventional Generation with Larger Sample Size ($n_{\mathcal{D_{\mathrm{test}}}}=1000$). We compare Cond-FiP against the baselines for the task of generating interventional data from the input noise variables. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario.Results indicate that Cond-FiP does not yet benefit from larger context sizes at inference, suggesting the need to scale both the model and training data for richer contexts.
K.2 Counterfactual Generation with Cond-FiP
We provide results (Table 30) for bechmarking Cond-FiP against baselines for the task of counterfactual generation. We operate in the same setup as the one in our main results ($n_{\mathcal{D_{\mathrm{test}}}}=400$Appendix C and all the methods are provided with the true casual graph. We observe that Unlike the tasks of noise prediction, sample & interventional generation, we find that Cond-FiP is worse than the baselines for the task of counterfactual generation. This can be explained as the training of Cond-FiP decoder relies on the true noise variables, and the model struggles to generalize the learned functional mechanisms when provided with inferred noise variables. We leave the improvement of Cond-FiP for counterfactual generation as future work.
Method Total Nodes LIN IN RFF IN LIN OUT RFF OUT
DoWhy 10 $0.03$ $(0.03)$ $0.13$ $(0.03)$ $0.0$ $(0.0)$ $0.04$ $(0.01)$
DECI 10 $0.1$ $(0.02)$ $0.2$ $(0.03)$ $0.04$ $(0.01)$ $0.11$ $(0.02)$
FiP 10 $0.03$ $(0.01)$ $0.09$ $(0.02)$ $0.02$ $(0.0)$ $0.03$ $(0.01)$
Cond-FiP 10 $0.09$ $(0.03)$ $0.21$ $(0.03)$ $0.05$ $(0.01)$ $0.11$ $(0.01)$
DoWhy 20 $0.01$ $(0.0)$ $0.12$ $(0.03)$ $0.0$ $(0.0)$ $0.13$ $(0.02)$
DECI 20 $0.06$ $(0.01)$ $0.15$ $(0.03)$ $0.07$ $(0.03)$ $0.15$ $(0.02)$
FiP 20 $0.03$ $(0.01)$ $0.1$ $(0.03)$ $0.06$ $(0.04)$ $0.09$ $(0.02)$
Cond-FiP 20 $0.09$ $(0.02)$ $0.26$ $(0.05)$ $0.13$ $(0.02)$ $0.3$ $(0.03)$
DoWhy 50 $0.0$ $(0.0)$ $0.09$ $(0.02)$ $0.0$ $(0.0)$ $0.17$ $(0.04)$
DECI 50 $0.04$ $(0.01)$ $0.11$ $(0.02)$ $0.03$ $(0.01)$ $0.18$ $(0.04)$
FiP 50 $0.03$ $(0.01)$ $0.08$ $(0.02)$ $0.03$ $(0.01)$ $0.14$ $(0.04)$
Cond-FiP 50 $0.1$ $(0.02)$ $0.26$ $(0.04)$ $0.1$ $(0.01)$ $0.46$ $(0.06)$
DoWhy 100 $0.0$ $(0.0)$ $0.08$ $(0.02)$ $0.0$ $(0.0)$ $0.2$ $(0.05)$
DECI 100 $0.02$ $(0.01)$ $0.1$ $(0.02)$ $0.02$ $(0.01)$ $0.22$ $(0.05)$
FiP 100 $0.01$ $(0.01)$ $0.07$ $(0.02)$ $0.02$ $(0.01)$ $0.19$ $(0.05)$
Cond-FiP 100 $0.09$ $(0.02)$ $0.29$ $(0.06)$ $0.13$ $(0.02)$ $0.56$ $(0.08)$
Table 30: Results for Counterfactual Generation. We compare Cond-FiP against the baselines for the task of generating counterfactual data from the input noise variables. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows denote the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results indicate that Cond-FiP struggles with counterfactual generation and cannot always match the performance of baselines trained from scratch.